# Explaining Datasets in Words: Statistical Models with Natural Language Parameters

### v1

@@ 6 Open-Ended Applications
We apply our framework to a broad range of applications to show that it is highly versatile. Our framework can monitor data streams (Section 6.1), apply to the visual domain (Section 6.2), and be easily steered to explain specific abstract properties (Section 6.3). Across all applications, our framework is able to explain complex concepts that classical methods struggle to produce.
@@ 6.1 Running Our Models Out of the Box: Monitoring Complex Data Streams of LLM Usages
We apply our models from Section 3.2 to monitor complex data streams of LLM usages. In particular, we recursively apply our clustering model to taxonomize user queries into application categories, apply our time series model to characterize trends in use cases across time, and apply our classification model to find categories where one LLM is better than the other. Due to space constraints, we present the key results in the main paper and the full results in Appendix F.
@@
To summarize, our framework 1) enables us to define a time series model to explain temporal trends in natural language, and 2) outputs sophisticated explanations that LDA fails to generate. However, it is far from perfect: it is slow to compute denotations for all pairs of $x$ and candidates $\phi$ since it involves many LLM API calls, and the predicates themselves are sometimes redundant. We describe these limitations and potential ways to improve them in Appendix F.
6.2 Applying Our Classification Model to Images: Explaining Memorable Visual Features
6.3 Explaining Abstract Properties via Easy Steering: Clustering Problems Based on Subarea
We apply our clustering model from Section 3.2 to cluster the MATH dataset [21] based on subareas. The MATH dataset contains five labeled subareas444after merging similar categories that differ in levels of difficulty and we hope our model can recover all of them: Algebra, counting__\__and__\__probability, geometry, number_theory, and precalculus. To steer our clustering model to explain subareas, we simply prompt the discretizer LLM “I want to cluster these math problems based on the type of skills required to solve them.” We set $K=5$, using gpt-4o to discretize and gpt-4o-mini to compute denotation.
@@ Appendix D Additional Results on Our Benchmark
To make sure that this takeaway is general and not specific to one embedding model, we run our method with another text embedding model, all-mpnet-base-v2555https://huggingface.co/blog/1b-sentence-embeddings and report the performance as the OtherEmb row. We find that using this neural embedding also outperforms OneHot in most cases, indicating that our conclusion is robust.
@@ Appendix E Detecting Self-Reinforcing Trends in Machine Learning System
Appendix F Implementation Details on Open-Ended Applications
@@
F.1 Taxonomizing User Applications.
@@
F.2 Characterizing Temporal Trends.
@@ F.3 Advantages and Limitations
Appendix G A Formal Description of Our Algorithm
@@
Appendix H Additional Details of Our paper
H.1 Limitations of Our Framework and Our Experiments
As mentioned in Appendix F.3, our current system is slow, as its performance highly depends on the LLMs to compute denotations correctly, it outputs semantically similar predicates that add little information, and it is hard to control the predicates to satisfy certain properties (e.g. being a subset of a parent category). Our experiments are limited since it assumes that the datasets and statistical models we used are reflective of real world application. We made our best effort to gather text clustering datasets that are commonly used in the literature (e.g. from Wang et al. [52], Pham et al. [38]) and defined models that are plausibly useful for practitioners. Additionally, note that our evaluation on topic clustering is more comprehensive than the prior work Wang et al. [52] by including two new datasets (Bills and Wiki); additionally, we used the exact same hyper-parameter across all clustering tasks, while Wang et al. [52] changed the hyper-parameters for different datasets.
H.2 Cost of the Experiments
@@
H.3 Licenses for Existing Datasets
@@
H.4 License for the Assets Provided by Our Paper
@@ 8. Answer: [Yes]
  9. Justification: The practical limitations are hinted at the end of Section 6 and more comprehensively discussed in Appendix F and H.1.
@@ 38. Answer: [Yes]
  39. Justification: We have described the cost of our experiments in Appendix H.2.
@@ 48. Answer: [Yes]
  49. Justification: We discussed them in Appendix H.5.
@@ 58. Answer: [Yes]
  59. Justification: We cited the existing datasets and mentioned their license in Appendix H.3.
@@ 63. Answer: [Yes]
  64. Justification: We discuss the license for the new assets introduced by our paper in Appendix H.4.

### v2

@@ 5.2 Metrics
For non-clustering tasks such as classification or time-series, we set the dimension of the learned predicates $\hat{\phi}(x)$ to be smaller than the reference predicates $\phi^{*}$, because some of the reference predicates are linearly dependent: for example, if there are four different languages in total (English, German, Spanish, and French), not being one of the first three implies that it is the last. Therefore, we set the dimension of the learned predicates $\hat{\phi}$ to be the number of linearly independent predicates in $\phi^{*}$. To compute the F1-score similarity, for each of the reference predicates, we find the most similar learned predicate, calculate the F1-score, and finally aggregate it across the reference predicates.
@@ 6 Open-Ended Applications
We apply our framework to a broad range of applications to show that it is highly versatile. Our framework can monitor data streams (Section 6.1), apply to the visual domain (Section F.1), and be easily steered to explain specific abstract properties (Section F.2). Across all applications, our framework is able to explain complex concepts that classical methods struggle to produce.
@@ 6.1 Running Our Models Out of the Box: Monitoring Complex Data Streams of LLM Usages
We apply our models from Section 3.2 to monitor complex data streams of LLM usages. In particular, we recursively apply our clustering model to taxonomize user queries into application categories, apply our time series model to characterize trends in use cases across time, and apply our classification model to find categories where one LLM is better than the other. Due to space constraints, we present the key results in the main paper and the full results in Appendix G.
@@
To summarize, our framework 1) enables us to define a time series model to explain temporal trends in natural language, and 2) outputs sophisticated explanations that LDA fails to generate. However, it is far from perfect: it is slow to compute denotations for all pairs of $x$ and candidates $\phi$ since it involves many LLM API calls, and the predicates themselves are sometimes redundant. We describe these limitations and potential ways to improve them in Appendix G.
Due to space constraints, we present applications in explaining visual features to make images memorable to humans and clustering math problems based on subareas in Appendix F.1 and F.2.
@@ Appendix D Additional Results on Our Benchmark
To make sure that this takeaway is general and not specific to one embedding model, we run our method with another text embedding model, all-mpnet-base-v2444https://huggingface.co/blog/1b-sentence-embeddings and report the performance as the OtherEmb row. We find that using this neural embedding also outperforms OneHot in most cases, indicating that our conclusion is robust.
@@ Appendix E Detecting Self-Reinforcing Trends in Machine Learning System
Appendix F More Applications
F.1 Applying Our Classification Model to Images: Explaining Memorable Visual Features
F.2 Explaining Abstract Properties via Easy Steering: Clustering Problems Based on Subarea
We apply our clustering model from Section 3.2 to cluster the MATH dataset [21] based on subareas. The MATH dataset contains five labeled subareas555after merging similar categories that differ in levels of difficulty and we hope our model can recover all of them: Algebra, counting__\__and__\__probability, geometry, number_theory, and precalculus. To steer our clustering model to explain subareas, we simply prompt the discretizer LLM “I want to cluster these math problems based on the type of skills required to solve them.” We set $K=5$, using gpt-4o to discretize and gpt-4o-mini to compute denotation.
Appendix G Implementation Details on Open-Ended Applications
@@
G.1 Taxonomizing User Applications.
@@
G.2 Characterizing Temporal Trends.
@@ G.3 Advantages and Limitations
Appendix H A Formal Description of Our Algorithm
@@
Appendix I Additional Details of Our paper
I.1 Limitations of Our Framework and Our Experiments
As mentioned in Appendix G.3, our current system is slow, as its performance highly depends on the LLMs to compute denotations correctly, it outputs semantically similar predicates that add little information, and it is hard to control the predicates to satisfy certain properties (e.g. being a subset of a parent category). Our experiments are limited since it assumes that the datasets and statistical models we used are reflective of real world application. We made our best effort to gather text clustering datasets that are commonly used in the literature (e.g. from Wang et al. [52], Pham et al. [38]) and defined models that are plausibly useful for practitioners. Additionally, note that our evaluation on topic clustering is more comprehensive than the prior work Wang et al. [52] by including two new datasets (Bills and Wiki); additionally, we used the exact same hyper-parameter across all clustering tasks, while Wang et al. [52] changed the hyper-parameters for different datasets.
I.2 Cost of the Experiments
@@
I.3 Licenses for Existing Datasets
@@
I.4 License for the Assets Provided by Our Paper
@@ 8. Answer: [Yes]
  9. Justification: The practical limitations are hinted at the end of Section 6 and more comprehensively discussed in Appendix G and I.1.
@@ 38. Answer: [Yes]
  39. Justification: We have described the cost of our experiments in Appendix I.2.
@@ 48. Answer: [Yes]
  49. Justification: We discussed them in Appendix I.5.
@@ 58. Answer: [Yes]
  59. Justification: We cited the existing datasets and mentioned their license in Appendix I.3.
@@ 63. Answer: [Yes]
  64. Justification: We discuss the license for the new assets introduced by our paper in Appendix I.4.